FINITE-VOLUME HYPERBOLIC 4-MANIFOLDS THAT SHARE A FUNDAMENTAL POLYHEDRON

DUBRAVKO IVANŠIĆ

Abstract. It is known that the volume function for hyperbolic manifolds of dimension ≥ 3 is finite-to-one. We show that the number of nonhomeomorphic hyperbolic 4-manifolds with the same volume can be made arbitrarily large. This is done by constructing a sequence of finite-sided finite-volume polyhedra with side-pairings that yield manifolds. In fact, we show that arbitrarily many nonhomeomorphic hyperbolic 4-manifolds may share a fundamental polyhedron. As a by-product of our examples, we also show in a constructive way that the set of volumes of hyperbolic 4-manifolds contains the set of even integral multiples of 4π^{2}/3. This is “half” the set of possible values for volumes, which is the integral multiples of 4π^{2}/3 due to the Gauss-Bonnet formula.